# Bayesian Sheaf Neural Networks

### v1

@@ 2.2 Sheaf Diffusion
In [6], sheaves with orthogonal restriction maps, rather than special orthogonal, are considered. However, since the sheaf learning mechanism in their SNN is a continuous parametrization of the orthogonal group $O(n)$ by the connected space $\mathbb{R}^{m}$, it must have its image contained in a connected component of $O(n)$. Given that this implies that the restriction maps $\mathcal{F}_{u\trianglelefteq e}$ will either all have determinant $1$ or all have determinant $-1$, the maps $\mathcal{F}_{u\trianglelefteq e}^{\mathsf{T}}\mathcal{F}_{v\trianglelefteq e}$ used to define the sheaf Laplacian must always have determinant equal to $1$. Thus it is effectively equivalent to learning restriction maps belonging to $SO(n)$.
@@
Proposition 2.3.
For a proof see Appendix A.
Remark 2.4.
An extra consequence of Proposition 2.3 is that it also provides a new upper bound on the stalk dimension $d$ necessary for $\mathcal{H}^{d}_{gen}$, the collection of sheaves with $d\times d$ general linear restriction maps, to have linear separation power over the collection of connected graph $\mathcal{G}_{C}$ with $C$ classes. This is due to the simple fact that $\mathcal{H}^{d}_{so}\subseteq\mathcal{H}^{d}_{gen}$.
3 Bayesian Sheaf Neural Networks
We briefly establish some notation. Let $P$ be a distribution on a measurable space $\mathcal{X}$ with density $p$ with respect to some reference measure $\mu$. If $X$ is a random variable with distribution $P$, we write $X\sim P$. For a function $f:\mathcal{X}\to\mathbb{R}$, we denote by $\mathbb{E}_{P}[f(X)]$ the expectation of $f(X)$ with respect to the distribution $P$, that is, $\mathbb{E}_{P}[f(X)]=\int_{\mathcal{X}}p(x)f(x)\,\mu(dx)$. If $Q$ is another distribution over $\mathcal{X}$, and $Q$ has density $q$ with respect to $\mu$, the Kullback-Leibler divergence of $P$ from $Q$ is
$\textnormal{KL}(P\;\|\;Q)=\int_{\mathcal{X}}p(x)\log\Bigg{(}\frac{p(x)}{q(x)}\Bigg{)}\,\mu(dx).$
To simplify notation, we will denote a distribution by its density function. For instance, an expression of the form $X\sim p(X|Y)$ denotes that the random variable $X$ is distributed according to the conditional distribution of $X$ given $Y$.
3.1 Problem Statement
$X$$A$$\phi$$\theta$$\mathcal{F}$$y$$L$
@@ 3.1 Problem Statement
Consider a semi-supervised node classification problem, in which we have a graph dataset $G=(V,E)$ together with a node feature matrix $X$, a collection of observed node labels, $y$, and a collection of unobserved node labels $y^{*}$, all of which are random samples. Motivated by the expressive power of the sheaf diffusion process, we model the conditional probability $p_{\theta}(y|X)$ using a sheaf neural network whose layers are given by Eq. 5, with network parameters $\theta$ and a cellular sheaf $\mathcal{F}$.
Rather than take the learned cellular sheaf $\mathcal{F}$ (or sheaves) to be a deterministic function of $X$ as in [6], we instead regard $\mathcal{F}$ as a latent random variable of our model. See Figure 1 for a depiction of the probability graphical model. Note that the posterior distribution $p_{\theta}(\mathcal{F}|X,y)$ will be intractable due to the fact that the labels $y$ are modeled as the output of a neural network. Hence, to do approximate posterior inference of $\mathcal{F}$ given the feature matrix, $X$, the observed node labels $y$, and model parameters $\theta$, we assume the intractable posterior distribution $p_{\theta}(\mathcal{F}|X,y)$ belongs to a predetermined family of variational distributions $q_{\phi}(\mathcal{F}|X,y)$ parameterized by weights $\phi$.
We begin by noting that the log-likelihood $\log p_{\theta}(X,y)$ can be written
@@
The quantity $\mathcal{L}(\theta,\phi;X,y)$ is known as the evidence lower bound or ELBO, due to the fact that the KL divergence from one distribution to another is always non-negative, hence $\log p_{\theta}(X,y)\geq\mathcal{L}(\theta,\phi,X,y)$. Note from Eq. (6) that maximizing the ELBO with respect to the parameters $\theta$ and $\phi$ has the effect of maximizing the log-likelihood of the data, as well as minimizing the KL divergence of the variational distribution $q_{\phi}(\mathcal{F}|X,y)$ from the true posterior distribution $p_{\theta}(\mathcal{F}|X,y)$. For this reason, the ELBO is commonly used as quantity to be optimized within variational Bayesian inference. See [5] for an overview of variational inference. In order to maximize $\mathcal{L}(\theta,\phi;X,y)$ with respect to the parameters $\theta,\phi$, we use the following equivalent expression:
@@
In many cases, a closed-form expression can be derived for the KL-divergence term in the right hand side of Eq. (8), leaving only the gradient of the expectation term with respect to $\theta$ and $\phi$ to be estimated. See Appendix C for calculations of the KL divergence in certain cases. Similar to [39], we estimate $\nabla_{\phi}\mathbb{E}_{q_{\phi}(\mathcal{F}|X,y)}[\log p_{\theta}(X,y|\mathcal{F})]$ by making use of the so-called reparameterization trick. In our case, for a random cellular sheaf $\mathcal{F}$ distributed as $\mathcal{F}\sim q_{\phi}(\mathcal{F}|X,y)$, the reparameterization trick expresses $\mathcal{F}$ as a deterministic function $\mathcal{F}=g_{\phi}(\epsilon,X)$ of random noise $\epsilon\sim p(\epsilon)$. This allows us to rewrite the expectation with respect to $q_{\phi}(\mathcal{F}|X,y)$ as an expectation with respect to $p(\epsilon)$ which in turn allows the gradient to be moved within the expectation. That is,
@@
~$\widetilde{\mathcal{L}}_{\lambda}(\theta,\phi;X,y)=\frac{1}{K}\sum_{k=1}^{K}\log p_{\theta}(X,y|\mathcal{F}^{(k)})-\lambda\,\textnormal{KL}(q_{\phi}(\mathcal{F}|X,y)\;\|\;p_{\theta}(\mathcal{F})).$ (9)
Here, the weight $\lambda$ of the KL divergence term is included to combat issues with the KL divergence term vanishing during training, which we cyclically anneal during training in a manner similar to [26]. When the KL divergence term of Eq. (8) cannot be computed analytically, we may estimate it in a similar manner by noting that $\textnormal{KL}(q_{\phi}(\mathcal{F}|X,y)\;\|\;p_{\theta}(\mathcal{F}))=\mathbb{E}_{q_{\phi}(\mathcal{F}|X,y)}[\log(q_{\phi}(\mathcal{F}|X,y)/p_{\theta}(\mathcal{F}))]$. In this case, the estimator, also denoted by ~$\widetilde{\mathcal{L}}_{\lambda}$, is given by
~$\widetilde{\mathcal{L}}_{\lambda}(\theta,\phi;X,y)=\frac{1}{K}\sum_{k=1}^{K}\Big{(}\log p_{\theta}(X,y|\mathcal{F}^{(k)})-\lambda\log\frac{q_{\phi}(\mathcal{F}^{(k)}|X,y)}{p_{\theta}(\mathcal{F}^{(k)})}\Big{)},$ (10)
@@ 3.2 Variational Sheaf Learner
We consider a variational sheaf learner in three different cases: where the learned sheaf $\mathcal{F}$ has restriction maps which are assumed to be (1) general linear, (2) special orthogonal, or (3) invertible diagonal. In all cases, we make the simplifying assumption that the variational distribution $q_{\phi}(\mathcal{F}|X,y)$ from the previous section factors as the product $\prod_{u\trianglelefteq e}q_{\phi}(\mathcal{F}_{u\trianglelefteq e}|X,y)$ over all incident node-edge pairs $u\trianglelefteq e$. While this assumption somewhat restricts the family of variational distributions, it makes computing the KL divergence term in the ELBO tractable. For further details, see Appendix C. We take each factor $q_{\phi}(\mathcal{F}_{u\trianglelefteq e}|X,y)$ to be a distribution $Q(\mu_{u\trianglelefteq e},\sigma_{u\trianglelefteq e})$ which depends on the type of restriction map used within the sheaf, where the distributional parameters $\mu_{u\trianglelefteq e},\sigma_{u\trianglelefteq e}$ are a function of the node features $X$ and $\phi$ as follows. For each incident node-edge pair $u\trianglelefteq e$, if $u^{\prime}\trianglelefteq e$ denotes the other node incident with $e$, a multi-layer perceptron with learned weights $\phi$, denoted $\operatorname{MLP}_{\phi}$, maps the concatenated vector $[x_{u}||x_{u^{\prime}}]$ to distributional parameters $\mu_{u\trianglelefteq e}$ and $\sigma_{u\trianglelefteq e}$:
$[\,\mu_{u\trianglelefteq e}\,||\,\sigma_{u\trianglelefteq e}\,]=\operatorname{MLP}_{\phi}([x_{u}||x_{u^{\prime}}]).$ (11)
@@
Case 1: General Linear. In the case where the sheaf restriction maps $\mathcal{F}_{u\trianglelefteq e}:\mathcal{F}(u)\to\mathcal{F}(e)$ are general linear maps, $\mu_{u\trianglelefteq e}$ and $\sigma_{u\trianglelefteq e}$ are $d^{2}$--dimensional vectors. We take $Q(\mu_{u\trianglelefteq e},\sigma_{u\trianglelefteq e})$ to be normal with diagonal covariance so that for $z\sim\mathcal{N}_{d^{2}}(\mu_{u\trianglelefteq e},\operatorname{diag}(\sigma_{u\trianglelefteq e}))$ we set $\mathcal{F}_{u\trianglelefteq e}$ to be the vector $z$ rearranged into a $d\times d$ matrix. The reparameterization trick in this case writes $z=\mu_{u\trianglelefteq e}+\operatorname{diag}(\sigma_{u\trianglelefteq e})\epsilon$ for $\epsilon\sim\mathcal{N}(\mathbf{0},I_{d^{2}})$ so that $\mathcal{F}_{u\trianglelefteq e}$, being the matrix obtained by reshaping the vector $z$, is a deterministic function of $\mu_{u\trianglelefteq e}$, $\sigma_{u\trianglelefteq e}$, and $\epsilon$. We use a standard normal prior, where $p_{\theta}(\mathcal{F})=\prod_{u\trianglelefteq e}p_{\theta}(\mathcal{F}_{u\trianglelefteq e})$ and $p_{\theta}(\mathcal{F}_{u\trianglelefteq e})=\mathcal{N}_{d^{2}}(\mathbf{0},I_{d^{2}})$.
Case 2: Special Orthogonal. The case where sheaf restriction maps are assumed to be special orthogonal presents challenges. Typical probability distribution on $SO(d)$, such as the matrix Langevin distribution, are not easily reparameterizable. This makes incorporating such distributions into neural networks, that are trained via back-propagation, difficult. On the other hand, while reparameterizable distributions on $SO(d)$ have been constructed in [21] using the exponential map, the density functions of such distributions can be difficult to compute exactly, which obstructs one from computing, or even efficiently estimating, the KL divergence term in Eq. (8). In Section 4 we define what we refer to as Cayley distributions, a family of reparameterizable distributions $\mathcal{C}_{d}(M,\kappa)$ on $SO(d)$ with closed form expressions for their probability density functions. We use $\mathcal{C}_{d}(M,\kappa)$ as the variational distributions for sheaves with special orthogonal restriction maps. Hence when $\mathcal{F}_{u\trianglelefteq e}\in SO(d)$, the distributional parameters are a mean rotation $\mu_{u\trianglelefteq e}\in SO(d)$ and scalar concentration parameter $\sigma_{u\trianglelefteq e}\in[0,1)$, and we have $Q(\mu_{u\trianglelefteq e},\sigma_{u\trianglelefteq e})=\mathcal{C}_{d}(\mu_{u\trianglelefteq e},\sigma_{u\trianglelefteq e})$. The reparameterization of the Cayley distributions is described in Remark 4.3. We use a prior distribution $p_{\theta}(\mathcal{F})=\prod_{u\trianglelefteq e}p_{\theta}(\mathcal{F}_{u\trianglelefteq e})$ where each $p_{\theta}(\mathcal{F}_{u\trianglelefteq e})$ is a uniform distribution on $SO(d)$.

### v2

@@ 2.2 Sheaf Diffusion
Although [10] consider sheaves with orthogonal restriction maps (i.e., elements of the orthogonal group $O(n)$), the sheaf learning mechanism in their SNN is a continuous parametrization of $O(n)$ by the connected space $\mathbb{R}^{m}$, so it must have its image contained in a connected component of $O(n)$. This implies that the restriction maps $\mathcal{F}_{u\trianglelefteq e}$ will either all have determinant $1$ or all have determinant 1-1. The maps $\mathcal{F}_{u\trianglelefteq e}^{\mathsf{T}}\mathcal{F}_{v\trianglelefteq e}$ used in the sheaf Laplacian must always have a determinant equal to $1$. Thus, it is effectively equivalent to learning restriction maps belonging to the special orthogonal group $SO(n)$.
2.3 Cayley Transform
Definition 7.
We denote the set of all $n\times n$ skew-symmetric matrices by $\mathfrak{so}(n)$. The Cayley transform[15] is a function $C:\mathfrak{so}(n)\to SO(n)$ defined by
$C(A)=(I-A)^{-1}(I+A)=2(I-A)^{-1}-I.$ (4)
If $A$ is an $n\times n$ skew-symmetric matrix, then $x^{\mathsf{T}}Ax=0$ for all $x\in\mathbb{R}^{n}$ due to the fact that $x^{\mathsf{T}}Ax=(x^{\mathsf{T}}Ax)^{\mathsf{T}}=-x^{\mathsf{T}}Ax$. Hence, if (IA)x=0(I-A)x=0, then $x^{\mathsf{T}}(I-A)x=x^{\mathsf{T}}x=0$ implies that $x=0$. Thus, IAI-A is invertible, showing that the Cayley transform is well-defined. If $P\in SO(n)$ does not have 1-1 as an eigenvalue, $C^{-1}(P)$ exists and is given by
$C^{-1}(P)=(P-I)(I+P)^{-1}=I-2(I+P)^{-1}.$ (5)
Remark 5.
For a vector $\phi\in\mathbb{R}^{n(n-1)/2}$, let $X_{\phi}$ denote the skew symmetric matrix containing the entries of $\phi$ as its lower triangular entries. Explicitly, if $\phi=(\phi_{1},\phi_{2},\dots,\phi_{n(n-1)/2})$ and if $x_{ij}$ denotes the $i,j$-th entry of $X_{\phi}$, then for $i>j$, set $x_{ij}=-x_{ji}=\phi_{ij-j(j+1)/2}$ and set $x_{ii}=0$. We may vectorize the Cayley transform and regard it as a map $\widetilde{C}:\mathbb{R}^{n(n-1)/2}\to\mathbb{R}^{n^{2}}$ by setting $\widetilde{C}(\phi)=\operatorname{vec}C(X_{\phi})$. The Jacobian (determinant) of $\widetilde{C}$ is $J\widetilde{C}(\phi)=\det(D\widetilde{C}(\phi)^{\mathsf{T}}D\widetilde{C}(\phi))^{1/2}$, where $D\widetilde{C}(\phi)$ denotes the total derivative of $\widetilde{C}$ at the point $\phi$. Following [50], the Jacobian $J\widetilde{C}(\phi)$ can be expressed in terms of $X_{\phi}$ by
$J\widetilde{C}(\phi)=\frac{2^{3n(n-1)/4}}{\det(I+X_{\phi})^{n-1}}.$ (6)
2.4 Variational Inference
Let $P$ be a distribution on a measurable space $\mathcal{X}$ with density $p$ with respect to some reference measure $\mu$. If $X$ is a random variable with distribution $P$, we write $X\sim P$. To simplify notation, we will often refer to a distribution by its density function.
For a function $f:\mathcal{X}\to\mathbb{R}$, we denote by $\mathbb{E}_{P}[f(X)]$ the expectation of $f(X)$ with respect to $P$, that is,
$\mathbb{E}_{P}[f(X)]=\int_{\mathcal{X}}p(x)f(x)\,\mu(dx).$ (7)
The Kullback-Leibler (KL) divergence is a fundamental measure from information theory that quantifies how one probability distribution diverges from a second, reference distribution [21]. If $Q$ is another distribution over $\mathcal{X}$ with density $q$ (also with respect to $\mu$), the KL divergence of $P$ from $Q$ is defined as
$\mathrm{KL}(P\,\|\,Q)=\int_{\mathcal{X}}p(x)\log\left(\frac{p(x)}{q(x)}\right)\mu(dx).$ (8)
Variational Inference (VI) transforms the problem of approximating intractable posterior distributions into an optimization task. Given a latent variable model with observations $y$ and latent variables $z$, the posterior $p(z|y)$ is often intractable due to the marginal likelihood requiring integration over the latent space. VI addresses this challenge by introducing a family of tractable distributions $q_{\phi}(z)$}, parameterized by variational parameters $\phi$, and finding the member of this family that is closest to the true posterior in terms of the KL divergence:
$\phi^{*}=\arg\min_{\phi}\operatorname{KL}\left(q_{\phi}(z)\|p(z\mid y)\right).$ (9)
This optimization is equivalent to maximizing the evidence lower bound (ELBO), which provides a lower bound on the marginal log-likelihood $\log p(y)$. The ELBO is given by
$\mathcal{L}(\phi)=\mathbb{E}_{q_{\phi}(z)}[\log p(y,z)]-\mathbb{E}_{q_{\phi}(z)}\left[\log q_{\phi}(z)\right].$ (10)
Intuitively, the ELBO balances two competing objectives. The first term encourages $q_{\phi}(z)$ to assign high probability to latent variables $z$ that explain the observed data $y$ well. The second term acts as a regularizer, penalizing complexity in the approximate posterior by keeping it close to the prior.
To see why the ELBO serves as a lower bound on $\log p(y)$, we note the identity:
$\log p(y)=\mathcal{L}(\phi)+\mathrm{KL}\left(q_{\phi}(z)\|p(z\mid y)\right),$ (11)
which shows that maximizing the ELBO is equivalent to minimizing the divergence from the true posterior. Since the KL divergence is always nonnegative, it follows that $\log p(y)\geq\mathcal{L}(\phi)$.
In practice, the approximate posterior $q_{\phi}(z)$ is often chosen from a parametric family such as multivariate Gaussians with diagonal covariance, which allows reparameterization techniques and gradient-based optimization. For a comprehensive treatment of variational inference and its theoretical foundations, see [6].
The choice of prior $p(z)$ plays an important role in variational inference. In many applications, a standard Gaussian prior $p(z)=\mathcal{N}(0,I)$ is used due to its simplicity and analytical tractability. Alternative priors, such as Laplace distributions for inducing sparsity or hierarchical priors for modeling structured uncertainty, can be used when specific inductive biases are desired.
2.5 Uncertainty Quantification Metrics
The two main types of uncertainty in uncertainty quantification are aleatoric uncertainty and epistemic uncertainty [24]. Aleatoric uncertainty arises from inherent data or environmental randomness, such as sensor noise or measurement variability, and is therefore irreducible. In contrast, epistemic uncertainty comes from limited data or an incomplete model and is reducible through increased data or refined modeling. In machine learning, aleatoric uncertainty manifests as output ambiguity, whereas epistemic uncertainty reflects model confidence given data limitations [42].
For a node classification task on a graph with $n$ nodes, $C$ classes, and an ensemble of $T$ stochastic forward passes, each input $x$ yields $T$ probability vectors $p_{t}(y|x)$, where $t=1,\ldots,T$. The ensemble predictive distribution is obtained by averaging the predictions: $\bar{p}(y\mid x)=\frac{1}{T}\sum_{t=1}^{T}p_{t}(y\mid x)$. The predicted class $\hat{y}$ is the class with the highest mean probability, given by $\hat{y}=\arg\max_{y}\bar{p}(y\mid x)$, and the associated confidence is $\hat{p}=\max_{y}\bar{p}(y\mid x)$. The variance across ensemble predictions for each class $c\in\{1,\ldots,C\}$ captures the model’s epistemic uncertainty, and is given by $\text{Var}_{c}(x)=\text{Var}_{t=1}^{T}\left[p_{t}(y=c\mid x)\right].$ Several uncertainty qualification metrics can be defined as follows.
Predictive entropy measures the total uncertainty in a model’s output (both aleatoric and epistemic). Lower entropy indicates that the model is more confident in its predictions [33].
$H[\bar{p}(y|x)]=-\sum_{c=1}^{C}\bar{p}(y=c\mid x)\log\bar{p}(y=c\mid x).$ (12)
Epistemic variance measures the variability in model predictions due to uncertainty in model parameters. Lower epistemic variance implies that the model’s predictions are consistent across stochastic forward passes [49].
$\text{EpistemicVariance}(x)=\frac{1}{C}\sum_{c=1}^{C}\text{Var}_{c}(x).$ (13)
Mutual information (MI) quantifies the epistemic component of uncertainty by measuring the divergence between the entropy of the mean prediction and the mean of the individual entropies. Lower MI values indicate reduced epistemic uncertainty, suggesting strong agreement among ensemble members and greater confidence in the model’s parameters [46].
$\text{MI}(x)=H\left[\bar{p}(y|x)\right]-\frac{1}{T}\sum_{t=1}^{T}H\left[p_{t}(y|x)\right].$ (14)
Expected Calibration Error (ECE) measures the discrepancy between predicted confidence scores and actual accuracy across a range of confidence intervals. To compute the ECE, the prediction confidences $\hat{p}_{i}$ are partitioned into $M$ equally spaced confidence intervals $I_{m}=\left(\frac{m-1}{M},\frac{m}{M}\right]$ for $m=1,\dots,M$. Let $B_{m}=\left\{i\in\{1,\dots,N\}\,\middle|\,\hat{p}_{i}\in I_{m}\right\}$ be the set of indices of samples whose confidence scores fall into interval $I_{m}$.
The accuracy of bin $B_{m}$ is defined as the proportion of correctly classified samples in that bin: $\text{acc}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\mathbbm{1}(\hat{y}_{i}=y_{i}),$ and the average confidence is $\text{conf}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\hat{p}_{i}.$
The Expected Calibration Error is computed as:
$\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{N}\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|.$ (15)
A lower ECE value indicates better calibration, meaning the predicted probabilities are more aligned with the observed accuracy [36].
3 Bayesian Sheaf Neural Networks
In this section, we introduce a Bayesian framework for Sheaf Neural Networks, extending the deterministic architecture by incorporating uncertainty in the learned sheaf structure.
3.1 Problem Statement
Consider a semi-supervised node classification problem, in which we have a graph dataset $G=(V,E)$ together with a node feature matrix $X$, a collection of observed node labels, $y$, and a collection of unobserved node labels, $y^{*}$, all of which are random samples. We model the conditional probability $p_{\theta}(y|X)$ using a sheaf neural network whose layers are given by Eq.(3), with network parameters $\theta$, adjacency matrix $A$, and a cellular sheaf $\mathcal{F}$.
Rather than taking the learned cellular sheaf $\mathcal{F}$ to be a deterministic function of $X$ , we regard $\mathcal{F}$ as a latent random variable of our model. See Figure 1 for a depiction of the probability graphical model. Note that the posterior distribution $p_{\theta}(\mathcal{F}|X,y)$ will be intractable due to the fact that the labels $y$ are modeled as the output of a neural network.
To approximate posterior inference of $\mathcal{F}$ given the feature matrix $X$, the observed node labels $y$, and model parameters $\theta$, we assume that the intractable posterior distribution $p_{\theta}(\mathcal{F}|X,y)$ belongs to a predetermined family of variational distributions $q_{\phi}(\mathcal{F}|X,y)$ parameterized by weights $\phi$ as discussed in Section 2.4. In our framework, we choose priors for $\mathcal{F}$ that reflect plausible variability in sheaf structures while preserving the tractability of inference.
@@ 3.1 Problem Statement
Following Eq.(11), the log-likelihood $\log p_{\theta}(X,y)$ can be written
@@
In order to maximize $\mathcal{L}(\theta,\phi;X,y)$ with respect to the parameters $\theta,\phi$, we use the following equivalent expression:
@@
In many cases, a closed-form expression can be derived for the KL divergence term in the right hand side of Eq.(18), leaving only the gradient of the expectation term with respect to $\theta$ and $\phi$ to be estimated. See Appendix Appendix C: KL Divergence Calculations for calculations of the KL divergence in certain cases.
Similar to [47], we estimate $\nabla_{\phi}\mathbb{E}_{q_{\phi}(\mathcal{F}|X,y)}[\log p_{\theta}(X,y|\mathcal{F})]$ by making use of the so-called reparameterization trick. In our case, for a random cellular sheaf $\mathcal{F}$ distributed as $\mathcal{F}\sim q_{\phi}(\mathcal{F}|X,y)$, the reparameterization trick expresses $\mathcal{F}$ as a deterministic function $\mathcal{F}=g_{\phi}(\epsilon,X)$ of random noise $\epsilon\sim p(\epsilon)$. This allows us to rewrite the expectation with respect to $q_{\phi}(\mathcal{F}|X,y)$ as an expectation with respect to $p(\epsilon)$ which in turn allows the gradient to be moved within the expectation. That is,
@@
$\widetilde{\mathcal{L}}_{\lambda}(\theta,\phi;X,y)=\frac{1}{K}\sum_{k=1}^{K}\log p_{\theta}(X,y|\mathcal{F}^{(k)})-\lambda\,\textnormal{KL}(q_{\phi}(\mathcal{F}|X,y)\;\|\;p_{\theta}(\mathcal{F})).$ (19)
Here, the weight $\lambda$ of the KL divergence term is included to combat issues with the KL divergence term vanishing during training, which we cyclically anneal during training in a manner similar to [31]. When the KL divergence term of Eq.(18) cannot be computed analytically, we may estimate it in a similar manner by noting that $\textnormal{KL}(q_{\phi}(\mathcal{F}|X,y)\;\|\;p_{\theta}(\mathcal{F}))=\mathbb{E}_{q_{\phi}(\mathcal{F}|X,y)}[\log(q_{\phi}(\mathcal{F}|X,y)/p_{\theta}(\mathcal{F}))]$. In this case, the estimator, also denoted by $\widetilde{\mathcal{L}}_{\lambda}$, is given by
$\widetilde{\mathcal{L}}_{\lambda}(\theta,\phi;X,y)=\frac{1}{K}\sum_{k=1}^{K}\Big{(}\log p_{\theta}(X,y|\mathcal{F}^{(k)})-\lambda\log\frac{q_{\phi}(\mathcal{F}^{(k)}|X,y)}{p_{\theta}(\mathcal{F}^{(k)})}\Big{)},$ (20)
@@ 3.2 Variational Sheaf Learner
We make the simplifying assumption that the variational distribution $q_{\phi}(\mathcal{F}|X,y)$ from the previous section factors as the product $\prod_{u\trianglelefteq e}q_{\phi}(\mathcal{F}_{u\trianglelefteq e}|X,y)$ over all incident node-edge pairs $u\trianglelefteq e$. While this assumption somewhat restricts the family of variational distributions, it makes computing the KL divergence term in the ELBO tractable. For further details, see Appendix Appendix C: KL Divergence Calculations.
We take each factor $q_{\phi}(\mathcal{F}_{u\trianglelefteq e}|X,y)$ to be a distribution $Q(\mu_{u\trianglelefteq e},\sigma_{u\trianglelefteq e})$ which depends on the type of restriction map used within the sheaf, where the distributional parameters $\mu_{u\trianglelefteq e},\sigma_{u\trianglelefteq e}$ are a function of the node features $X$ and $\phi$ as follows. For each incident node-edge pair $u\trianglelefteq e$, if $u^{\prime}\trianglelefteq e$ denotes the other node incident with $e$, a multi-layer perceptron with learned weights $\phi$, denoted $\operatorname{MLP}_{\phi}$, maps the concatenated vector $x_{u}||x_{u^{\prime}}]$ to distributional parameters $\mu_{u\trianglelefteq e}$ and $\sigma_{u\trianglelefteq e}$:
$\mu_{u\trianglelefteq e}\,||\,\sigma_{u\trianglelefteq e}\,]=\operatorname{MLP}_{\phi}([x_{u}||x_{u^{\prime}}]).$ (21)
@@
The distribution $Q$ for the variational sheaf learners is considered in three settings, where the learned sheaf $\mathcal{F}$ has restriction maps assumed to be: (1) general linear, (2) invertible diagonal, or (3) special orthogonal.
@@
Case 1: General Linear. In the case where the sheaf restriction maps $\mathcal{F}_{u\trianglelefteq e}:\mathcal{F}(u)\to\mathcal{F}(e)$ are general linear maps, $\mu_{u\trianglelefteq e}$ and $\sigma_{u\trianglelefteq e}$ are $d^{2}$-dimensional vectors. We take $Q(\mu_{u\trianglelefteq e},\sigma_{u\trianglelefteq e})$ to be normal with diagonal covariance so that for $z\sim\mathcal{N}_{d^{2}}(\mu_{u\trianglelefteq e},\operatorname{diag}(\sigma_{u\trianglelefteq e}))$ we set $\mathcal{F}_{u\trianglelefteq e}$ to be the vector $z$ rearranged into a $d\times d$ matrix. The reparameterization trick in this case writes $z=\mu_{u\trianglelefteq e}+\operatorname{diag}(\sigma_{u\trianglelefteq e})\epsilon$ for $\epsilon\sim\mathcal{N}(\mathbf{0},I_{d^{2}})$ so that $\mathcal{F}_{u\trianglelefteq e}$, being the matrix obtained by reshaping the vector $z$, is a deterministic function of $\mu_{u\trianglelefteq e}$, $\sigma_{u\trianglelefteq e}$, and $\epsilon$. We use a standard normal prior, where $p_{\theta}(\mathcal{F})=\prod_{u\trianglelefteq e}p_{\theta}(\mathcal{F}_{u\trianglelefteq e})$ and $p_{\theta}(\mathcal{F}_{u\trianglelefteq e})=\mathcal{N}_{d^{2}}(\mathbf{0},I_{d^{2}})$.